### v1

# MCVO: A Generic Visual Odometry for Arbitrarily Arranged Multi-Cameras

Abstract
Making multi-camera visual SLAM systems easier to set up and more robust to the environment is always one of the focuses of vision robots. Existing monocular and binocular vision SLAM systems have narrow FoV and are fragile in textureless environments with degenerated accuracy and limited robustness. Thus multi-camera SLAM systems are gaining attention because they can provide redundancy for texture degeneration with wide FoV. However, current multi-camera SLAM systems face massive data processing pressure and elaborately designed camera configurations, leading to estimation failures for arbitrarily arranged multi-camera systems. To address these problems, we propose a generic visual odometry for arbitrarily arranged multi-cameras, which can achieve metric-scale state estimation with high flexibility in the cameras’ arrangement. Specifically, we first design a learning-based feature extraction and tracking framework to shift the pressure of CPU processing of multiple video streams. Then we use the rigid constraints between cameras to estimate the metric scale poses for robust SLAM system initialization. Finally, we fuse the features of the multi-cameras in the SLAM back-end to achieve robust pose estimation and online scale optimization. Additionally, multi-camera features help improve the loop detection for pose graph optimization. Experiments on KITTI-360 and MultiCamData datasets validate the robustness of our method over arbitrarily placed cameras. Compared with other stereo and multi-camera visual SLAM systems, our method obtains higher pose estimation accuracy with better generalization ability. Our codes and online demos are available at https://github.com/JunhaoWang615/MCVO.
I Introduction
Visual Simultaneous Localization and Mapping (vSLAM) is a fundamental technique in robotics and autonomous navigation, enabling a system to estimate its motion relative to its environment using visual sensors. Traditional monocular or stereo SLAM systems often require strict camera configurations and rely heavily on integrating inertial measurement units (IMUs) for true scale pose estimation and map construction [1, 2]. However, these approaches are limited by the narrow FoV and perform poorly in environments where camera placement flexibility is paramount, such as most cars with 6 different cameras. Recently, multi-camera visual SLAM allows for greater robustness in these situations, which can cover more surrounding visual scenes to provide redundancy for poorly textured environments. It opens up new possibilities for applications such as UAV navigating through cluttered environments [3], and autonomous driving with surrounding multi-cameras [4].
Figure 1: Illustration of the proposed MCVO system. An example of using two fisheye and one pinhole camera on KITTI-360 dataset [5] for state estimation. The proposed MCVO obtains better performance compared with ORB-SLAM3 [2] using the front stereo camera.
However, current multi-camera visual SLAM systems still face challenges for real-world applications. On the one hand, the increase in the number of cameras, while providing information redundancy, inevitably brings several times the pressure of data processing. Most existing methods adopt conventional feature association methods, such as ORB [6, 4], which leads to a sharp increase in CPU usage and difficulty in balancing the resources for the back nonlinear optimization, resulting in the requirement of high-performance CPU or failure to achieve real-time. On the other hand, accurate scale estimation is also challenging for MCVO systems. Existing methods mostly estimate the scale by multiple binocular setups or adding IMUs [7, 8, 9], which require elaborate configuration of the camera’s FoV overlap or extrinsic calibration between camera and IMU sensors, making it difficult to balance the flexibility of the system configuration and the accuracy of the scale estimation. Therefore, the main objective of this work is to achieve robust and generalized multi-camera visual odometry by solving the feature association computation pressure and scale estimation problem for arbitrarily arranged multi-camera systems.
To address these challenges, we propose a generic multi-camera visual odometry system, i.e., MCVO, which takes only multiple rigidly bundled cameras at arbitrary orientations and positions. We first analyze the state-of-the-art feature association methods from the accuracy-efficiency perspective and design a learning-based feature extraction and tracking framework to shift the computation pressure of CPU processing of multiple video streams. Then, we initialize the SLAM system to obtain true scale body poses based on the rigid constraint between the aligned poses for each camera using SfM. In the backend, we fuse multi-camera features to achieve robust pose estimation and scale optimization. The multi-camera features are further concatenated in Bag Of Words (BoW) for loop closure detection. Through rigorous testing and practical implementation on the KITTI-360 and MultiCamData datasets, we aim to demonstrate the efficacy of our system in enhancing the capabilities of autonomous agents across various domains with unprecedented levels of flexibility and generalizability. The highlights of the proposed system: (i) It enables the use of multiple cameras positioned in any orientation with only the requirement of the extrinsic parameters. (ii) By eliminating the dependency on IMU and providing the flexibility to handle cameras in arbitrary configurations, it focuses solely on visual information, no matter overlapping or non-overlapping cameras, which allows for true scale estimation and online optimization, leading to enhanced accuracy and robustness. (iii) Besides, it can accommodate a variety of camera types, including but not limited to fisheye and standard pinhole cameras, making it suitable for a broad range of applications. The main contributions are listed as follows:
  • We propose a generic visual SLAM system framework for arbitrarily arranged multi-cameras, with novel designs for front-end multi-camera feature detection and matching, system initialization, back-end optimization, and loop closure.
  • We comprehensively consider the existing SOTA feature association methods and design a SuperPoint-based extractor and LK tracking front-end, shifting the computational pressure of the CPU and improving the stability of feature matching, which finally provides a guarantee for the stability and efficiency of multi-eye SLAM.
  • We propose a SLAM scale estimation strategy based on the trajectory consistency of multi-cameras, which is compatible with different camera models and with or without camera FoV overlap, demonstrating strong generalization and stability.
II Related Work
II-A Efficient Feature Association Frontend
[…]
To enhance system efficiency, some approaches utilize GPUs for parallel processing, shifting the computationally intensive front-end of multi-camera systems to reduce CPU usage and latency. Common feature detectors, such as Shi-Tomasi and Harris, have been implemented in CUDA Visual Library (VILIB) [11] for fast feature extraction. He et al. [9] developed front-end feature tracking algorithms that run on GPUs using NVIDIA’s VPI. Additionally, learning-based methods for feature extraction and tracking [12][13] have been proposed to enhance feature robustness while reducing the computational burden on CPU. Pandey et al. [14] introduced a visual odometry algorithm combining deep learning with optical flow tracking.
To further reduce CPU usage, only selecting useful features for backend optimization can improve pose estimation accuracy while lowering the computational demands. For example, RANSAC [15] is a common practice to remove obvious outliers with large reprojection errors. Scoring on the feature descriptor and uncertainty estimation using networks are also often used for selecting features for SLAM systems [16, 17]. Carlone et al. [18] introduced an efficient feature selection algorithm using the Max-logDet metric and minimal eigenvalue, which Zhao et al. [19] later improved by optimizing the feature subset selection to maximize logDet, which makes the algorithm an order of magnitude faster than state-of-the-arts. These feature selection strategies are useful for reducing the number of features for multi-camera visual odometry systems.
II-B Multiple Camera Visual Odomotry
Most current VO algorithms focus on monocular and stereo cameras, such as VINS-MONO [1] and ORB-SLAM [2]. While these methods offer good performance and efficiency, they are still not robust enough for autonomous pose estimation in complex real-world environments. Multi-camera setups can significantly increase visual constraints and provide richer image information, enhancing the robustness of SLAM systems. As a result, multi-camera SLAM has gained growing attention. Early work by Sola et al. [20] explored multi-camera VO by fusing multiple monocular camera data using filtering methods.
Figure 2: The pipeline of the proposed MCVO system. Scale estimation and correction ensure scale stability.
More recent multi-camera SLAM systems [3, 6] extend monocular PTAM [21] and ORB-SLAM to synchronized multi-camera rigs through optimization-based approaches. Liu et al. [4] proposed a pose tracker and local mapper capable of supporting arbitrary numbers of stereo cameras. Tribou et al. [22] introduced a real-time visual odometry system utilizing multiple non-overlapping FoV cameras. This configuration provides a wider FoV, enhancing redundancy for backend optimization. However, it is challenging for non-overlapping cameras to obtain accurate scale information. To improve scale accuracy, some works adopt multi-camera setups with overlapping FoV [23][24]. Zhao et al. [25] proposed an early multi-camera depth estimation method, achieving stable long-term depth estimation. Kaveti et al. [7] developed a multi-camera framework that extracts 3D feature points and real scale by matching overlapping areas between cameras. Xu et al. [26] proposed a depth estimation method that does not require large FoV overlaps, enhancing depth estimation generalization across different multi-camera configurations. In addition to traditional scale estimation methods, many deep learning-based approaches have emerged [27][28]. While these methods perform well on specific datasets, their heavy reliance on contextual information results in poor generalization in cross-dataset experiments [29].
While overlapping multi-camera setups improve the scale estimation accuracy, they do so by reducing the range of observed information, which limits the applicability of multi-camera SLAM in non-overlapping configurations. To a certain extent, this also goes against the original intention of using multi-camera systems to maximize FoV and information capture. Our proposed MCVO accurately estimates scale information based on multi-camera trajectory consistency, regardless of camera overlaps. It can be applied to arbitrarily arranged camera setups, fully leveraging the wide FoV advantages of multi-camera systems.
III Methodology
III-A Overview
Our proposed generic multi-camera visual odometry framework is illustrated in Fig. 2. The main inputs to the framework are the synchronized multi-camera video sequences. The multiple cameras are rigidly bundled and calibrated in advance with known intrinsic and extrinsic parameters. The output is the metric scale 6-DoF robot poses in the real-world environment. The pipeline comprises four components: frontend feature extraction, pose and metric scale initialization, backend optimization, and loop closure. To accelerate the frontend of multi-camera feature association, we employ feature extraction with GPU acceleration and 3-priority feature selection (Section III-B. Then we initialize the multi-camera system of metric scale body poses using the multi-camera poses and extrinsic parameters (Section III-C). To ensure the realism of the motion scale, we perform an adaptive correlation of the scale deviations during the backend optimization (Section III-D). Given the larger FoV of the multi-camera system, we design a more robust multi-camera omnidirectional loop detection algorithm (Section III-E). We further optimize the body poses in the loop with the pose graph constraint.
III-B Efficient Multi-Camera Frontend
The frontend of our proposed MCVO primarily processes multi-camera images, extracts and matches useful features, and then provides reliable landmarks for the backend optimization.
Improving feature detection quality and association efficiency are essential for the visual odometry frontend. As the number of cameras increases, the volume of extracted and tracked features grows significantly, leading to a considerable computational burden for real-time performance. Regarding these issues, we utilize SuperPoint [30] for feature extraction and LK optical flow for feature tracking. This process is accelerated via GPU, substantially reducing CPU usage brought by multiple cameras and enhancing performance in handling image streams.
We further conduct a feature selection process, to reduce the CPU load during optimization and improve estimation performance. The traditional quad-tree algorithm [31] tends to place new features close to existing features in texture-rich regions, which might reduce the optimization accuracy due to insufficient constraints. Our feature selection algorithm avoids excessive feature concentration to ensure an even distribution of features through three key priorities.
  1. 1.
    Quantity priority: Prioritizes nodes with more features to enhance processing efficiency.
  2. 2.
    Tracked feature priority: Prioritizes reducing the density of points in texture-rich areas by avoiding adding useless features near tracked points.
  3. 3.
    High score feature priority: Prioritizes features with higher scores.
Figure 3: Distribution of feature points. The proposed approach achieves more uniform feature distribution and better robustness than the score-based method. Red points: tracked features. Blue points: newly added features.
III-C Multi-Camera Initialization
The initialization of our MCVO system aims to estimate initial poses and determine the metric scale of motions. This process includes the pose initialization for each camera and metric scale estimation for body poses.
We first calculate the parallax for each camera in a 10-frame window. When all the camera frames at time $t$ have sufficient parallax (greater than 30 pixels), we set the world coordinate origin at time $t$ for all cameras. The camera with the most tracking features at time $t$ is selected as the principal camera. Simultaneously, we estimate the poses for each camera independently using Structure from Motion (SfM) on the landmarks provided by the frontend for the 10 frames in the initialization window. It should be noted that the SfM is conducted on 10 frames of each monocular camera, therefore, the scale of each camera is not on the same metric level.
Subsequently, given the poses for all cameras, we first calculate the body frame (principal camera frame) poses using the camera extrinsic parameter
$\bm{p}^{b_{c}}_{t}=\bm{R}^{c}_{t}\bm{p}^{c}+s^{c}\bm{p}^{c}_{t},$ (1)
where $\bm{R}^{c}_{t}$, $\bm{p}^{c}_{t}$ are the estimated rotation and translation for camera $c$ at timestamp $t$, $s^{c}$ is the scale parameter for camera $c$ to be estimated, $\bm{p}^{c}$ is the translational extrinsic parameter of camera $c$ and $\bm{P}^{b_{c}}_{t}$ is the body frame translation estimated from camera $c$ at timestamp $t$.
However, because each camera’s pose is computed independently, their scales are not aligned, resulting in significant discrepancies in body poses. Theoretically, the estimated body pose should be consistent across different cameras. Therefore, we take the differences between body poses calculated by different cameras as the residuals and define the optimization problem as follows:
$\min_{s_{1},s_{2},...,s_{n_{c}}}\sum_{t}\sum_{i=1}^{n_{c}}\sum_{j=i+1}^{n_{c}}$ $||\bm{p}^{b_{i}}_{t}-\bm{p}^{b_{j}}_{t}||^{2},$ (2)
where $n_{c}$ is the number of cameras, $s_{k}(k\in[1,n_{c}])$ is the scale parameter for camera $k$, $\bm{P}^{b_{c}}_{t}$ is the body frame translation estimated from camera $k$ at timestamp $t$. To solve the least squares problem for Eq. 2, we use the Levenberg-Marquardt algorithm from the Ceres Solver library [32].
We readjust camera poses according to the estimated metric scale, then re-triangulate all features using the updated poses to obtain their metric depths. After the initial estimation of poses and scales, we proceed with a sliding window-based backend optimization for state estimation.
III-D Scale Correction and Backend Optimization
After initializing the pose estimation system with metric scale. We use a vision-only bundle adjustment with the extracted features from all cameras to minimizes the sum of prior and the Mahalanobis norm of all visual measurement residuals to obtain a maximum posterior state estimation,
$\min_{\mathcal{X}}$$\Big{\{}\|\bm{r}_{0}\|^{2}_{\mathbf{\Sigma}_{0}}+\sum_{i\in\mathcal{C}}\sum_{j\in\mathcal{F}_{i}}{\bm{\rho}(\|\bm{r}_{v_{(i,j)}}\|}^{2}_{\mathbf{\Sigma}_{{v}_{(i,j)}}})\Big{\}}$ (3)
$\mathcal{X}$ $=[\bm{x}_{1}^{b},\bm{x}_{2}^{b},...,\bm{x}_{t}^{b},\bm{\lambda}_{1},\bm{\lambda}_{2},...,\bm{\lambda}_{n_{c}}]$
where the state vector $\mathcal{X}$ includes the body poses $\bm{x}_{t}^{b}$ in the sliding window and all the feature depths $\bm{\lambda}_{c}$ observed by the $c$-th camera, $\bm{\rho}(\cdot)$ is the robust huber loss, $\mathcal{C}$ is the set of cameras, $\mathcal{F}_{i}$ is the set of reliable features in $i$-th camera, $\bm{r}_{0}$ is the marginalization factor, $\bm{r}_{v_{(i,j)}}$ is the $j$-th landmark reprojection residual in the $i$-th camera. The definition of all the residuals can be found in [1]. The Dogleg algorithm in the Ceres Solver [32] is employed to solve this least squares problem.
Scale Correction. The optimization in Eq. 3 continuously provides accurate body poses and feature depths. However, scale error gradually accumulates as the map expands. To address this, we conduct scale correction alongside the backend optimization at regular intervals. During scale correction, a vision-only bundle adjustment is first performed for each camera to determine its poses independently of the body. We then apply the scale estimation algorithm described in Section III-C to obtain the scale factors $\bm{s}$. Instead of directly adjusting the pose scales, we modify the feature depths in each camera, allowing for a smoother scale correction of poses through subsequent state estimation.
III-E Multi-Camera Loop Closure
Although the multi-camera VO and scale correction provide high-accuracy state estimation, the accumulated drifts still exist alongside the running of pure visual odometry. To mitigate the drifts over time, we design a multi-camera loop closure module across different cameras and establish loop connections in the pose graph for further state optimization.
Loop Detection. In our MCVO pipeline, the loop detection simply leverages DBoW2 [33] algorithm for place recognition. The key point is how to use multi-camera features to improve the loop detection accuracy and robustness. Firstly, the features from all cameras are stored in the corresponding bag-of-words model. Then potential loop frames are identified by comparing matching scores between DBoW2 models from different frames. The wide field of view offered by multi-camera setups improve the detection rate of potential loops.
Figure 4: Illustration of the multi-camera re-localization
Upon detecting a loop, we proceed with re-localization as illustrated in Fig. 4. The current frame features are converted into 3D features $\bm{z}$ with their depth information. By matching feature descriptors from the looped frames, we establish 2D-3D feature correspondences $\{(\bm{x}_{\textit{i}},\bm{z}_{\textit{i}})\}$. When a camera has sufficient 2D-3D feature pairs, we proceed with the following optimization to refine camera poses
$\min_{\mathcal{R},\mathcal{P}}\sum_{\mathcal{C}}\sum_{i}\big{|}\big{|}\bm{x}_{\textit{i}}-\bm{\pi}(\bm{K}_{\textit{c}},\bm{z}_{\textit{i}},\bm{R},\bm{p})\big{|}\big{|}^{2},$ (4)
where $\bm{K}_{\textit{c}}$ is the intrinsic and the extrinsic (to body frame) parameters of camera $c$, $\bm{\pi}$ is the projection of the $\bm{z}_{\textit{i}}$ on the image plane. The least squares problem is also solved by the Levenberg-Marquardt algorithm.
This process ensures accurate re-localization even with significant viewpoint differences. If multiple cameras meet the criteria, we use extrinsic constraints to verify the relocalization’s success further. Once verified, the accurate body pose of the loop frame is determined based on the camera poses and extrinsic parameters. Compared to a monocular system, our multi-camera setup enhances loop detection from different viewpoints, significantly increasing the detection rate of loops.
TABLE I: Comparison of ATE on various datasets. The metrics translation, rotation, and scale are in meters, degrees, and percentages, respectively. Note that MultiCamSLAM fails after running a certain distance (as shown in parentheses) on the KITTI360.
Method KITTI360_00 KITTI360_03 KITTI360_05 KITTI360_10 MCdata_Lab1 MCdata_Ground1
(1520m) (1379m) (1173m) (3343m) (152m) (90m)
Trans. Rot. Scal. Trans. Rot. Scal. Trans. Rot. Scal. Trans. Rot. Scal. Trans. Trans. Trans. Trans.
VINS-Fusion 76.198 114.923 1.003 38.213 125.079 0.585 14.749 123.580 0.181 135.536 130.248 2.924 16.579 10.921% 14.80 16.469%
ORBSLAM3 22.673 117.981 0.390 12.269 115.832 0.306 8.134 114.361 0.164 44.317 114.268 1.484 2.90 1.91% 0.85 0.80%
MultiCamSLAM 15.68(245.2) 115.516 0.341 8.91(222.5) 114.082 0.184 11.28(91.9) 155.944 0.150 56.46(553.5) 136.175 1.561 14.83 9.75% 1.69 1.88%
MCVO(ORB) 44.054 8.634 0.518 64.861 4.803 1.367 49.143 5.353 0.591 102.050 6.139 2.833 11.50 7.561% 14.58 16.321%
MCVO(ours) 13.933 6.847 0.300 20.234 2.585 0.394 6.895 3.341 0.079 76.282 7.348 2.638 7.894 5.200% 1.54 1.711%
Figure 5: Comparison of ATE for different methods on the KITTI360
Pose Graph Optimization We follow the commonly used 4-DoF pose graph optimization [1] to update the trajectory in the loops. The pose graph $\bm{G}=\{\bm{V},\bm{E}\}$ keep track of all keyframes with their spatial constraints, where vertices $\bm{V}$ represent keyframes and edges $\bm{E}$ store the transformation constraints between keyframe poses. Keyframes marginalized during the sliding window are added to the pose graph, with their relative transformations to other keyframes in the local sliding window incorporated as edges. When a loop is detected, a loop edge between the current frame and the loop frame is added to $\bm{E}$, followed by the 4-DoF pose graph optimization. The optimization process effectively reduces accumulated errors and enhances the system’s robustness.
IV Experiments
In this section, we conduct experiments on two public datasets to evaluate our proposed method.
IV-A Experimental Setup
IV-A1 Datasets
For the evaluation of our method, the dataset must contain multi-camera setup with pre-calibrated intrinsic and extrinsic parameters. We select two public available dataset, i.e., KITTI360[5] and MultiCamData[7], which have different camera types, scenario types, and various camera layouts. The KITTI360 dataset was collected in outdoor road environments with two $180^{\circ}$ fisheye cameras and two $90^{\circ}$ pinhole cameras, as shown in Fig. 1. The four cameras are synchronized and pre-calibrated. Compared to other datasets such as EuRoC [34], KITTI360 features larger scale, longer distances, and challenging scenarios, such as roads under bridges (lighting changes), wilderness (texture degradation), and dynamic scenes. We tested on most of the sequences and compared with the ground truth trajectories. MultiCamData was collected in indoor environments using 6 cameras. We mainly use the non-overlapping cameras (cam0, cam5, and cam6) for generalizability. The dataset also includes challenging scenarios, such as narrow corridors, large white walls (texture degradation), sudden turns, and dynamic scenes.
IV-A2 Evaluation Setup
We select VINS-Fusion [1], ORB-SLAM3 [2], and MultiCamSLAM [7] as baselines for comparison. To obtain metric scale state estimation for VINS-Fusion and ORB-SLAM3, these two methods are setup with front stereo cameras of each dataset. While for MultiCamSLAM, we employ a non-overlapping camera setup consistent with the method [7]. MCVO is configured with two fisheye cameras and one front pinhole camera on KITTI360 dataset. The evaluation metrics on KITTI360 include Absolute Trajectory Error (ATE), Relative Pose Error (RPE) and scale error [35]. Since MultiCamData lacks ground truth trajectory, we estimate the errors between the initial and final poses using a visual target as in [7], given that the start and end of the trajectories are at the same location. All the experiments are conducted on a desktop with a 2.5GHz Intel Core i7-11700 CPU and an NVIDIA RTX3060 GPU.
IV-B Experimental Results
The quantitative results for ATE are presented in Table I and Fig. 5. Compared with the stereo ORB-SLAM3 and VINS-Fusion, MCVO achieves significantly lower rotation error and competitive translation error on most of the sequences of KITTI360, primarily due to the wide FoV provided by the multi-camera setup and the high accuracy and robustness of the frontend. The small scale errors further demonstrate the effectiveness of our scale estimation strategy. The MultiCamSLAM method cannot finish most of the sequences on KITTI360. This issue arises because the KITTI360 dataset was captured at high speeds, leading to large inter-frame displacements. MultiCamSLAM struggles with significant inter-frame displacements, hindering triangulation and initialization. The comparison with MCVO on ORB features further shows the robustness of our front-end design. Besides, MCVO shows strong performance and robustness on MultiCamData. It outperforms VINS-Fusion in accuracy, with a marginal shortfall compared to ORB-SLAM3. This slight reduction in accuracy is primarily attributed to the presence of numerous white walls in indoor corridors, which introduce noise that impacts multi-camera VO. Compared to MultiCamSLAM, which also employs a non-overlapping camera setup, MCVO achieves superior accuracy and better generalization across different sequences.
Fig. 6 presents the RPE results of different methods on the 00 sequence of KITTI360 dataset. MCVO significantly outperforms other methods with smallest pose drift. This improvement is largely due to the wide coverage of scenes in the multi-camera setup with a robust and efficient frontend. The expanded FoV offers stronger constraints for state estimation, resulting in higher pose estimation accuracy. Additionally, the robust frontend ensures accurate feature tracking even in low-quality scenes. These results validate the effectiveness of MCVO on state estimation.
Figure 6: Comparison of RPE for different methods on the KITTI360_00 sequence.
To further qualitative analyze the performance, we plot the trajectories of different methods on sequences 00 and 05 of KITTI360 dataset in Fig. 7. ORB-SLAM3 and VINS-Fusion exhibit large rotation errors on the 00 sequence. In contrast, MCVO shows improved stability and accuracy under such conditions. Additionally, the overall trajectory of our approach is smoother with more stable position estimation.
Figure 7: Trajectories of different sequences on KITTI360.
IV-C Ablation Study
In this section, we conduct ablation experiments for different numbers of cameras, feature extractors, feature matching, and whether to use loop closure. The effectiveness of each part of our proposed method is verified through comprehensive analyses.
TABLE II: MCVO performance under different number of cameras on the KITTI360 00 sequence.
Method Transl.[m] Rot.[°] UsageCPU[%]
Mean Std Mean Std Mean
2 cams 45.523 43.042 8.180 7.263 203.2
3 cams 13.933 13.640 6.847 6.806 316.7
4 cams 20.673 22.262 7.671 7.812 443.6
IV-C1 Camera number
Table II shows the state estimation accuracy of MCVO under different number of cameras. MCVO on two cameras obtains relatively lower accuracy due to the limited FoV, which provides insufficient information. When the number of cameras increases to 3, the estimation accuracy reaches its peak. However, with 4 cameras, there is a slight decrease in accuracy. This is primarily because two fisheye and one perspective cameras have already covered enough surrounding information, the added new camera will bring more observation noises and computational burden for the system. Consequently, the gain from the new camera cannot contribute to improving accuracy.
TABLE III: MCVO performance under different feature extractors on the KITTI360 00 sequence.
Method Transl.[m] Rot.[°] UsageCPU[%] FPS
Mean Std Mean Std Mean
Shi-Tomasi 39.390 40.612 9.256 8.290 398.3 18.75
ORB 44.054 39.900 8.634 8.327 942.4 12.74
SuperPoint 13.933 13.640 6.847 6.806 316.7 19.50
IV-C2 Feature extractor
We use pose accuracy and CPU usage metrics to compare different feature extraction methods. Table III shows that SuperPoint achieves the highest accuracy compared with commonly used Shi-Tomasi and ORB. This is mainly because SuperPoint can extract multi-level features and is more robust to environmental changes. Additionally, the CPU usage of MCVO(SuperPoint) is significantly lower. This reduction in CPU usage is because of our GPU-accelerated frontend. Consequently, The shift of CPU usage also ensures the computation resource for back-end optimization of accurate pose estimation and efficiency.
IV-C3 Feature matching
Table IV shows the performance of MCVO on different feature-matching methods. Flow-based methods exhibit lower CPU usage and more inliers compared to matching-based methods (KNN, SuperGlue). Additionally, we count the mean tracking times of each feature. LK shows the longest tracking times compared to others, which provides stable landmarks for the backend pose estimation. The small number of track times of other methods leads to the pose initialization failures.
TABLE IV: MCVO performance under different feature matching methods (Superglue-SG).
Method Inlier UsageCPU Aver. track Initialization
Num ratio Mean times succ.
ORB+KNN 51.33 34.22% 892.03% 2.71
ORB+LK 93.04 62.03% 750.78% 15.95
SuperPoint+SG 65.23 43.39% 789.53% 4.32
SuperPoint+LK 99.61 66.41% 710.12% 16.40
IV-C4 Loop closure
We also conduct experiments on the KITTI360 05 sequence to validate the effectiveness of the loop closure. The qualitative results are shown in Fig. 8. We can observe that the multi-camera loop closure effectively performs loop detection and closure, providing stable global constraints for back-end optimization. MCVO successfully corrects the cumulative error in the map, achieving more precise trajectory estimation.
Figure 8: Loop closure effectiveness validation on the KITTI360 05 sequence.
IV-D Discussion
The comparison with other stereo VO and multi-camera VO demonstrates the effectiveness of the proposed MCVO on state estimation. The analyses of each component validate the reasonableness of our design on multi-camera data processing. However, there are still limitations:
  • Multi-camera configurations can improve VO robustness, but inevitably introduce new computational requirements. Our approach uses learning-based feature extraction to shift CPU consumption but introduces the need for GPU resources. The best approach may be to develop a specialized camera with feature association embedded in the sensor hardware.
  • MCVO requires the extrinsic parameters of arbitrarily arranged multi-cameras. When there is no overlap between FoV, the extrinsic calibration will be a challenging problem.
V Conclusions and Future Work
In this paper, we propose a generic multi-camera visual odometry system, which has high flexibility on the camera setup with different camera types, with or without FoV overlap. The arbitrary arrangement of multi-cameras, not only improves the flexibility of the SLAM system but also achieves robust camera-only true-scale pose estimation. The designed learning-based feature association front-end utilizes GPU to effectively release the computational pressure of CPUs on multi-camera data processing and improves the stability of feature matching. The system functionalities of multi-camera features for scale initialization, back-end optimization, and loop closure are fully validated during the experiments. Comparisons with the state-of-the-art further demonstrate the effectiveness and robustness of MCVO on state estimation. Future work will focus on the deployment of MCVO on more platforms.

### v2

@@
 Abstract
-Making multi-camera visual SLAM systems easier to set up and more robust to the environment is always one of the focuses of vision robots.
-Existing monocular and binocular vision SLAM systems have narrow FoV and are fragile in textureless environments with degenerated accuracy and limited robustness.
-Thus multi-camera SLAM systems are gaining attention because they can provide redundancy for texture degeneration with wide FoV.
-However, current multi-camera SLAM systems face massive data processing pressure and elaborately designed camera configurations, leading to estimation failures for arbitrarily arranged multi-camera systems.
-To address these problems, we propose a generic visual odometry for arbitrarily arranged multi-cameras, which can achieve metric-scale state estimation with high flexibility in the cameras’ arrangement.
-Specifically, we first design a learning-based feature extraction and tracking framework to shift the pressure of CPU processing of multiple video streams.
-Then we use the rigid constraints between cameras to estimate the metric scale poses for robust SLAM system initialization.
-Finally, we fuse the features of the multi-cameras in the SLAM back-end to achieve robust pose estimation and online scale optimization.
+Making multi-camera visual SLAM systems easier to set up and more robust to the environment is attractive for vision robots.
+Existing monocular and binocular vision SLAM systems have narrow sensing Field-of-View (FoV), resulting in degenerated accuracy and limited robustness in textureless environments.
+Thus multi-camera SLAM systems are gaining attention because they can provide redundancy with much wider FoV.
+However, the usual arbitrary placement and orientation of multiple cameras make the pose scale estimation and system updating challenging.
+To address these problems, we propose a robust visual odometry system for rigidly-bundled arbitrarily-arranged multi-cameras, namely MCVO, which can achieve metric-scale state estimation with high flexibility in the cameras’ arrangement.
+Specifically, we first design a learning-based feature tracking framework to shift the pressure of CPU processing of multiple video streams to GPU.
+Then we initialize the odometry system with the metric-scale poses under the rigid constraints between moving cameras.
+Finally, we fuse the features of the multi-cameras in the back-end to achieve robust pose estimation and online scale optimization.
 Additionally, multi-camera features help improve the loop detection for pose graph optimization.
-Experiments on KITTI-360 and MultiCamData datasets validate the robustness of our method over arbitrarily placed cameras.
-Compared with other stereo and multi-camera visual SLAM systems, our method obtains higher pose estimation accuracy with better generalization ability.
-Our codes and online demos are available at https://github.com/JunhaoWang615/MCVO.
+Experiments on KITTI-360 and MultiCamData datasets validate its robustness over arbitrarily arranged cameras.
+Compared with other stereo and multi-camera visual SLAM systems, our method obtains higher pose accuracy with better generalization ability.
+The code will also be made publicly available soon.
 I Introduction
 Visual Simultaneous Localization and Mapping (vSLAM) is a fundamental technique in robotics and autonomous navigation, enabling a system to estimate its motion relative to its environment using visual sensors.
-Traditional monocular or stereo SLAM systems often require strict camera configurations and rely heavily on integrating inertial measurement units (IMUs) for true scale pose estimation and map construction [1, 2].
-However, these approaches are limited by the narrow FoV and perform poorly in environments where camera placement flexibility is paramount, such as most cars with 6 different cameras.
+Traditional monocular or stereo SLAM systems often require strict camera configurations and rely heavily on integrating inertial measurement units (IMUs) for metric-scale pose estimation and map construction [1, 2].
+However, these approaches are limited by the narrow FoV and perform poorly in environments where camera placement flexibility is paramount, such as most cars with 6 cameras.
 Recently, multi-camera visual SLAM allows for greater robustness in these situations, which can cover more surrounding visual scenes to provide redundancy for poorly textured environments.
@@
 Figure 1: Illustration of the proposed MCVO system.
-An example of using two fisheye and one pinhole camera on KITTI-360 dataset [5] for state estimation.
-The proposed MCVO obtains better performance compared with ORB-SLAM3 [2] using the front stereo camera.
+An example of using 3-cam and 4-cam setups on KITTI-360 dataset [5] for state estimation.
+The proposed MCVO performs better than ORB-SLAM3 [2] using the front stereo camera.
 However, current multi-camera visual SLAM systems still face challenges for real-world applications.
@@
 Therefore, the main objective of this work is to achieve robust and generalized multi-camera visual odometry by solving the feature association computation pressure and scale estimation problem for arbitrarily arranged multi-camera systems.
-To address these challenges, we propose a generic multi-camera visual odometry system, i.e., MCVO, which takes only multiple rigidly bundled cameras at arbitrary orientations and positions.
-We first analyze the state-of-the-art feature association methods from the accuracy-efficiency perspective and design a learning-based feature extraction and tracking framework to shift the computation pressure of CPU processing of multiple video streams.
+To address these challenges, we propose a robust multi-camera visual odometry system, i.e., MCVO, which takes only multiple rigidly bundled cameras at arbitrary orientations and positions, outputting high-precision metric-scale body poses.
+We first design a learning-based feature extraction and tracking framework to shift the computation pressure of CPU to GPU.
 Then, we initialize the SLAM system to obtain true scale body poses based on the rigid constraint between the aligned poses for each camera using SfM.
@@
 The multi-camera features are further concatenated in Bag Of Words (BoW) for loop closure detection.
-Through rigorous testing and practical implementation on the KITTI-360 and MultiCamData datasets, we aim to demonstrate the efficacy of our system in enhancing the capabilities of autonomous agents across various domains with unprecedented levels of flexibility and generalizability.
+Through rigorous testing and practical implementation on the KITTI-360 and MultiCamData datasets, we aim to demonstrate the effectiveness of our MCVO system in enhancing the capabilities of autonomous agents across various domains with unprecedented levels of flexibility and generalizability.
 The highlights of the proposed system: (i) It enables the use of multiple cameras positioned in any orientation with only the requirement of the extrinsic parameters.
@@
 The main contributions are listed as follows:
-  • We propose a generic visual SLAM system framework for arbitrarily arranged multi-cameras, with novel designs for front-end multi-camera feature detection and matching, system initialization, back-end optimization, and loop closure.
-  • We comprehensively consider the existing SOTA feature association methods and design a SuperPoint-based extractor and LK tracking front-end, shifting the computational pressure of the CPU and improving the stability of feature matching, which finally provides a guarantee for the stability and efficiency of multi-eye SLAM.
-  • We propose a SLAM scale estimation strategy based on the trajectory consistency of multi-cameras, which is compatible with different camera models and with or without camera FoV overlap, demonstrating strong generalization and stability.
+  • We propose a generic visual SLAM system framework based on universal multi-camera topology, which innovatively incorporates multi-camera feature extraction frontend, metric-scale pose initialization and optimization, and multi-camera loop closure.
+  • We pioneer a robust odometry scale estimation strategy for arbitrarily arranged multi-cameras.
+Through establishing multi-camera manifold consistency constraints on motion trajectories, we effectively calculate metric-scale poses for multi-camera systems in the initialization.
+Subsequently, we continuously refine metric scale in the backend optimization.
+  • We develop a cascaded tracking algorithm integrating SuperPoint features and Lucas-Kanade optical flow, effectively reducing computational load while enhancing loop closure through multi-camera collaborative verification.
 II Related Work
@@
 [9] developed front-end feature tracking algorithms that run on GPUs using NVIDIA’s VPI.
-Additionally, learning-based methods for feature extraction and tracking [12][13] have been proposed to enhance feature robustness while reducing the computational burden on CPU.
+Additionally, learning-based methods for feature extraction and tracking [12][13] have been proposed to enhance feature robustness while reducing the computational burden on the CPU.
 Pandey et al.
@@
 These feature selection strategies are useful for reducing the number of features for multi-camera visual odometry systems.
-II-B Multiple Camera Visual Odomotry
+II-B Multiple Camera Visual Odometry
 Most current VO algorithms focus on monocular and stereo cameras, such as VINS-MONO [1] and ORB-SLAM [2].
@@
 III-A Overview
-Our proposed generic multi-camera visual odometry framework is illustrated in Fig.
+Our proposed multi-camera visual odometry framework is illustrated in Fig.
 2.
@@
 The pipeline comprises four components: frontend feature extraction, pose and metric scale initialization, backend optimization, and loop closure.
-To accelerate the frontend of multi-camera feature association, we employ feature extraction with GPU acceleration and 3-priority feature selection (Section III-B.
-Then we initialize the multi-camera system of metric scale body poses using the multi-camera poses and extrinsic parameters (Section III-C).
-To ensure the realism of the motion scale, we perform an adaptive correlation of the scale deviations during the backend optimization (Section III-D).
-Given the larger FoV of the multi-camera system, we design a more robust multi-camera omnidirectional loop detection algorithm (Section III-E).
+To accelerate the frontend of multi-camera feature association, we employ feature extraction with GPU acceleration and 3-priority feature selection.
+Then we initialize the multi-camera system with metric-scale poses using the motion invariance constraints and extrinsic parameters.
+To ensure the realism of the motion scale, we perform an adaptive correlation of the scale deviations during the backend optimization.
+Given the larger FoV of the multi-camera system, we design a more robust multi-camera omnidirectional loop detection algorithm.
 We further optimize the body poses in the loop with the pose graph constraint.
 III-B Efficient Multi-Camera Frontend
-The frontend of our proposed MCVO primarily processes multi-camera images, extracts and matches useful features, and then provides reliable landmarks for the backend optimization.
-Improving feature detection quality and association efficiency are essential for the visual odometry frontend.
-As the number of cameras increases, the volume of extracted and tracked features grows significantly, leading to a considerable computational burden for real-time performance.
-Regarding these issues, we utilize SuperPoint [30] for feature extraction and LK optical flow for feature tracking.
-This process is accelerated via GPU, substantially reducing CPU usage brought by multiple cameras and enhancing performance in handling image streams.
-We further conduct a feature selection process, to reduce the CPU load during optimization and improve estimation performance.
-The traditional quad-tree algorithm [31] tends to place new features close to existing features in texture-rich regions, which might reduce the optimization accuracy due to insufficient constraints.
-Our feature selection algorithm avoids excessive feature concentration to ensure an even distribution of features through three key priorities.
+The frontend of our proposed MCVO primarily processes multi-camera images, extracts and matches useful features, and then provides reliable landmarks for system initialization and backend optimization.
+Enhancing feature detection quality and association efficiency is critical for the visual odometry frontend.
+With an increasing number of cameras, the volume of extracted and tracked features grows exponentially, imposing a substantial computational burden on the edge-computing device.
+To address these challenges, we implement SuperPoint [30] for feature extraction and LK optical flow for tracking.
+Leveraging GPU acceleration, this framework significantly reduces CPU overhead from multi-camera setups while improving image stream processing efficiency.
+To mitigate CPU load during optimization and enhance estimation accuracy, we propose a feature selection algorithm, the 3-priority quad-tree, which ensures uniform feature distribution.
+The conventional quadtree method clusters feature in high-texture regions, resulting in insufficient geometric constraints and compromised backend optimization accuracy.
+In contrast, our approach enforces spatial diversity through three criteria to mitigate feature over-concentration.
   1. 1.
-    Quantity priority: Prioritizes nodes with more features to enhance processing efficiency.
+    Quantity priority: Prioritize nodes with more features to enhance processing efficiency.
   2. 2.
-    Tracked feature priority: Prioritizes reducing the density of points in texture-rich areas by avoiding adding useless features near tracked points.
+    Tracked feature priority: Prioritize reducing the density of points in texture-rich areas by avoiding adding useless features near tracked points.
   3. 3.
-    High score feature priority: Prioritizes features with higher scores.
+    High score feature priority: Prioritize features with higher scores.
 Figure 3: Distribution of feature points.
@@
 III-C Multi-Camera Initialization
-The initialization of our MCVO system aims to estimate initial poses and determine the metric scale of motions.
-This process includes the pose initialization for each camera and metric scale estimation for body poses.
-We first calculate the parallax for each camera in a 10-frame window.
-When all the camera frames at time $t$ have sufficient parallax (greater than 30 pixels), we set the world coordinate origin at time $t$ for all cameras.
-The camera with the most tracking features at time $t$ is selected as the principal camera.
-Simultaneously, we estimate the poses for each camera independently using Structure from Motion (SfM) on the landmarks provided by the frontend for the 10 frames in the initialization window.
-It should be noted that the SfM is conducted on 10 frames of each monocular camera, therefore, the scale of each camera is not on the same metric level.
-Subsequently, given the poses for all cameras, we first calculate the body frame (principal camera frame) poses using the camera extrinsic parameter
-$\bm{p}^{b_{c}}_{t}=\bm{R}^{c}_{t}\bm{p}^{c}+s^{c}\bm{p}^{c}_{t},$ (1)
-where $\bm{R}^{c}_{t}$, $\bm{p}^{c}_{t}$ are the estimated rotation and translation for camera $c$ at timestamp $t$, $s^{c}$ is the scale parameter for camera $c$ to be estimated, $\bm{p}^{c}$ is the translational extrinsic parameter of camera $c$ and $\bm{P}^{b_{c}}_{t}$ is the body frame translation estimated from camera $c$ at timestamp $t$.
-However, because each camera’s pose is computed independently, their scales are not aligned, resulting in significant discrepancies in body poses.
-Theoretically, the estimated body pose should be consistent across different cameras.
-Therefore, we take the differences between body poses calculated by different cameras as the residuals and define the optimization problem as follows:
-$\min_{s_{1},s_{2},...,s_{n_{c}}}\sum_{t}\sum_{i=1}^{n_{c}}\sum_{j=i+1}^{n_{c}}$ $||\bm{p}^{b_{i}}_{t}-\bm{p}^{b_{j}}_{t}||^{2},$ (2)
-where $n_{c}$ is the number of cameras, $s_{k}(k\in[1,n_{c}])$ is the scale parameter for camera $k$, $\bm{P}^{b_{c}}_{t}$ is the body frame translation estimated from camera $k$ at timestamp $t$.
-To solve the least squares problem for Eq.
-2, we use the Levenberg-Marquardt algorithm from the Ceres Solver library [32].
-We readjust camera poses according to the estimated metric scale, then re-triangulate all features using the updated poses to obtain their metric depths.
-After the initial estimation of poses and scales, we proceed with a sliding window-based backend optimization for state estimation.
+The initialization phase of our MCVO system addresses two critical objectives: establishing initial camera poses and recovering the metric scale of motion trajectories.
+This process involves two components: camera pose initialization and cross-sensor metric scale calibration.
+Our approach initiates by analyzing feature parallaxes within a 10-frame sliding window across all cameras.
+When the parallax of all camera streams exceeds 30 pixels at timestamp $t$, we establish the world coordinate system anchored at this timestamp.
+The camera exhibiting the highest feature tracking stability at $t$ is designated as the principal camera.
+Concurrent with this spatial anchoring, we perform independent Structure-from-Motion (SfM) reconstructions for each camera using their respective observed landmarks within the initialization window.
+The pose estimations for each camera $c$ in the sliding window are mathematically represented as $\mathcal{R}^{c}$ and $\mathcal{T}^{c}$, where $\bm{R}^{c}_{t}\in\text{SO}(3)$ denotes the rotation matrix and $\bm{T}^{c}_{t}\in\mathbb{R}^{3\times 1}$ represents the translation vector at frame $t$.
+Then we have
+$\mathcal{R}^{c}$ =$=\left\{\bm{R}^{c}_{t},t\in[0,10]\right\},$ (1)
+$\mathcal{T}^{c}$ =$=\left\{\bm{T}^{c}_{t},t\in[0,10]\right\}.$
+It should be emphasized that these monocular SfM instances yield scale-ambiguous poses, necessitating subsequent metric scale unification through multi-camera constraints.
+Concretely, building upon the camera pose estimations, we can derive the body frame poses (attached to the principal camera) through extrinsic calibration parameters $\bm{r}^{c}$ and $\bm{t}^{c}$.
+Notably, the scale ambiguity inherent in monocular SfM causes inter-camera scale misalignment, manifesting as inconsistent body pose estimates across sensor streams.
+To resolve this scale discrepancy, we introduce scale factors $\bm{s}^{c}\in\mathbb{R}^{+}$ for each camera to enable metric-scale body pose estimation:
+$\begin{bmatrix}\bm{R}^{b_{c}}_{t}&\bm{T}^{b_{c}}_{t}\\ \bm{0}&\bm{1}\end{bmatrix}$ $=\begin{bmatrix}\bm{R}^{c}_{t}&s^{c}\bm{T}^{c}_{t}\\ \bm{0}&\bm{1}\end{bmatrix}\begin{bmatrix}(\bm{r}^{c})^{T}&-(\bm{r}^{c})^{T}\bm{t}^{c}\\ \bm{0}&\bm{1}\end{bmatrix}$ (2)
+$=\begin{bmatrix}\bm{R}^{c}_{t}\bm{(}{r}^{c})^{T}&-\bm{R}^{c}_{t}(\bm{r}^{c})^{T}\bm{t}^{c}+s^{c}\bm{T}^{c}_{t}\\ \bm{0}&\bm{1}\end{bmatrix}$
+Leveraging the theoretical motion consistency constraint that body trajectories should be scale-invariant across camera perspectives, we formulate a residual term for camera $i$ and $j$:
+$\bm{\hat{e}}(\bm{T}^{b_{i}}_{t},\bm{T}^{b_{j}}_{t})=$ $\bm{T}^{b_{i}}_{t}-\bm{T}^{b_{j}}_{t}$ (3)
+$=$ $-\bm{R}^{i}_{t}(\bm{r}^{i})^{T}\bm{t}^{i}+s^{i}\bm{T}^{i}_{t}+\bm{R}^{j}_{t}(\bm{r}^{j})^{T}\bm{t}^{j}-s^{j}\bm{T}^{j}_{t}$
+$=$ $\begin{bmatrix}\bm{T}^{i}_{t}&-\bm{T}^{j}_{t}\end{bmatrix}\begin{bmatrix}s^{i}\\ s^{j}\end{bmatrix}+\bm{\theta},$
+where $\theta=-\bm{R}^{i}_{t}(\bm{r}^{i})^{T}\bm{t}^{i}+\bm{R}^{j}_{t}(\bm{r}^{j})^{T}\bm{t}^{j}$.
+Since camera rotation matrices $\bm{R}^{c}_{t}$ are scale-independent, $\bm{\theta}$ is related to the camera extrinsic parameters $[\bm{r}^{c},\bm{t}^{c}]$ and camera rotation matrices $\bm{R}^{c}_{t}$, thus $\bm{\theta}$ will be also scale-independent.
+Building upon the pairwise residual formulation in (3), we extend the constraint to $N$-camera systems through block matrix construction.
+Let the scale vector encapsulate all camera scale factors, the residual takes the form:
+$\bm{e}(\bm{T}^{b_{i}}_{t},\bm{T}^{b_{j}}_{t})$ $=\begin{bmatrix}\bm{0}_{\alpha}&\bm{T}^{i}_{t}&\bm{0}_{\beta}&-\bm{T}^{j}_{t}&\bm{0}_{\gamma}\end{bmatrix}\bm{s}+\bm{\theta}$ (4)
+$=\bm{F}\bm{s}+\bm{\theta},$
+$\bm{s}$ $=\begin{bmatrix}s^{1}&s^{2}&\ldots&s^{N}\end{bmatrix}^{T}$
+where $\alpha=3\times(i-1)$, $\beta=3\times(j-i-1)$, $\gamma=3\times(N-j),(i<j)$.
+$\bm{F}\in\mathbb{R}^{3\times N}$ forms the sparse observation matrix.
+This construction preserves the problem’s sparsity pattern for efficient large-scale optimization.
+$\bm{\theta}$ denotes the $\bm{s}$-independent component in Equation (3).
+The complete scale estimation problem then reduces to solving the constrained least-squares minimization:
+$\min_{s_{1},s_{2},...,s_{N}}\sum_{t}\sum_{i=1}^{N}\sum_{j=i+1}^{N}$ $||\bm{e}(\bm{T}^{b_{i}}_{t},\bm{T}^{b_{j}}_{t})||^{2}$ (5)
+The scale estimation is formulated as a manifold-aware optimization problem within the ceres solver [31], with scale factor $\bm{s}$ parameterized over Lie groups.
+We solve the optimization problem using the Levenberg-Marquardt algorithm, which effectively handles the sparse Jacobian structure, with automatic differentiation ensuring numerical precision in partial derivative calculations.
 III-D Scale Correction and Backend Optimization
-After initializing the pose estimation system with metric scale.
-We use a vision-only bundle adjustment with the extracted features from all cameras to minimizes the sum of prior and the Mahalanobis norm of all visual measurement residuals to obtain a maximum posterior state estimation,
-$\min_{\mathcal{X}}$$\Big{\{}\|\bm{r}_{0}\|^{2}_{\mathbf{\Sigma}_{0}}+\sum_{i\in\mathcal{C}}\sum_{j\in\mathcal{F}_{i}}{\bm{\rho}(\|\bm{r}_{v_{(i,j)}}\|}^{2}_{\mathbf{\Sigma}_{{v}_{(i,j)}}})\Big{\}}$ (3)
-$\mathcal{X}$ $=[\bm{x}_{1}^{b},\bm{x}_{2}^{b},...,\bm{x}_{t}^{b},\bm{\lambda}_{1},\bm{\lambda}_{2},...,\bm{\lambda}_{n_{c}}]$
-where the state vector $\mathcal{X}$ includes the body poses $\bm{x}_{t}^{b}$ in the sliding window and all the feature depths $\bm{\lambda}_{c}$ observed by the $c$-th camera, $\bm{\rho}(\cdot)$ is the robust huber loss, $\mathcal{C}$ is the set of cameras, $\mathcal{F}_{i}$ is the set of reliable features in $i$-th camera, $\bm{r}_{0}$ is the marginalization factor, $\bm{r}_{v_{(i,j)}}$ is the $j$-th landmark reprojection residual in the $i$-th camera.
-The definition of all the residuals can be found in [1].
-The Dogleg algorithm in the Ceres Solver [32] is employed to solve this least squares problem.
+Following metric scale initialization, we formulate a multi-camera visual bundle adjustment problem that jointly optimizes camera poses and 3D landmarks by minimizing the Mahalanobis-weighted reprojection residuals:
+$\min_{\mathcal{X}}$$\Big{\{}\|\bm{r}_{0}\|^{2}_{\mathbf{\Sigma}_{0}}+\sum_{c\in N}\sum_{f\in\mathcal{F}_{c}}{\bm{\rho}(\|\bm{r}_{v_{(c,f)}}\|}^{2}_{\mathbf{\Sigma}_{{v}_{(c,f)}}})\Big{\}},$ (6)
+$\mathcal{X}$ $=[\bm{x}_{1}^{b},\bm{x}_{2}^{b},...,\bm{x}_{t}^{b},\bm{\lambda}_{1},\bm{\lambda}_{2},\cdots,\bm{\lambda}_{N}],$
+where $\bm{x}_{t}^{b}$ denotes the body poses in the sliding window and $\bm{\lambda}_{c}$ denotes inverse-depth parameterized landmarks observed in camera $c$.
+$\mathcal{F}_{i}$ is the set of reliable features in $i$-th camera, $\bm{r}_{0}$ is the marginalization factor, $\bm{r}_{v_{(i,j)}}$ is the $j$-th landmark reprojection residual in the $i$-th camera.
+$\bm{\rho}(\cdot)$ implements Huber robustification.
+The Ceres solver’s Dogleg algorithm [31] efficiently handles this nonlinear least-squares problem through adaptive trust region management and analytical Jacobians derived from [1].
 Scale Correction.
-The optimization in Eq.
-3 continuously provides accurate body poses and feature depths.
-However, scale error gradually accumulates as the map expands.
-To address this, we conduct scale correction alongside the backend optimization at regular intervals.
-During scale correction, a vision-only bundle adjustment is first performed for each camera to determine its poses independently of the body.
-We then apply the scale estimation algorithm described in Section III-C to obtain the scale factors $\bm{s}$.
-Instead of directly adjusting the pose scales, we modify the feature depths in each camera, allowing for a smoother scale correction of poses through subsequent state estimation.
+While the iterative bundle adjustment process continuously refines body poses and feature depths through nonlinear optimization, inherent scale ambiguity leads to progressive scale drift during map expansion.
+To mitigate this critical issue, we implement a hierarchical scale correction module that operates during the backend optimization.
+Subsequently, we compute the optimal scale factors $s$ through the constrained optimization framework detailed in Section III-C.
+Rather than applying direct scale transformation to pose trajectories, we optimally adjust inverse feature depths based on scale factors within each camera’s coordinate system.
+This differential depth propagation approach enables implicit scale rectification through subsequent state estimation updates while maintaining the temporal consistency of the SLAM system.
 III-E Multi-Camera Loop Closure
-Although the multi-camera VO and scale correction provide high-accuracy state estimation, the accumulated drifts still exist alongside the running of pure visual odometry.
-To mitigate the drifts over time, we design a multi-camera loop closure module across different cameras and establish loop connections in the pose graph for further state optimization.
+Despite high-accuracy multi-camera VO and scale correction, drift accumulation persists. To address the problem, we design a cross-camera loop closure module.
 Loop Detection.
-In our MCVO pipeline, the loop detection simply leverages DBoW2 [33] algorithm for place recognition.
-The key point is how to use multi-camera features to improve the loop detection accuracy and robustness.
-Firstly, the features from all cameras are stored in the corresponding bag-of-words model.
-Then potential loop frames are identified by comparing matching scores between DBoW2 models from different frames.
-The wide field of view offered by multi-camera setups improve the detection rate of potential loops.
+We construct a cross-camera feature fusion mechanism based on the DBow2[32] model.
+By integrating synchronously captured features from multi-camera into a unified feature space and enforcing spatiotemporal consistency constraints for inter-frame feature matching, we establish cross-camera shared bag-of-words representations.
+Specifically, as shown in Fig.
+4.
+This methodology effectively mitigates geometric constraint degradation caused by viewpoint disparities, significantly enhancing the robustness of loop detection in complex multi-camera environments.
 Figure 4: Illustration of the multi-camera re-localization
-Upon detecting a loop, we proceed with re-localization as illustrated in Fig.
-4.
-The current frame features are converted into 3D features $\bm{z}$ with their depth information.
-By matching feature descriptors from the looped frames, we establish 2D-3D feature correspondences $\{(\bm{x}_{\textit{i}},\bm{z}_{\textit{i}})\}$.
-When a camera has sufficient 2D-3D feature pairs, we proceed with the following optimization to refine camera poses
-$\min_{\mathcal{R},\mathcal{P}}\sum_{\mathcal{C}}\sum_{i}\big{|}\big{|}\bm{x}_{\textit{i}}-\bm{\pi}(\bm{K}_{\textit{c}},\bm{z}_{\textit{i}},\bm{R},\bm{p})\big{|}\big{|}^{2},$ (4)
-where $\bm{K}_{\textit{c}}$ is the intrinsic and the extrinsic (to body frame) parameters of camera $c$, $\bm{\pi}$ is the projection of the $\bm{z}_{\textit{i}}$ on the image plane.
-The least squares problem is also solved by the Levenberg-Marquardt algorithm.
-This process ensures accurate re-localization even with significant viewpoint differences.
-If multiple cameras meet the criteria, we use extrinsic constraints to verify the relocalization’s success further.
-Once verified, the accurate body pose of the loop frame is determined based on the camera poses and extrinsic parameters.
-Compared to a monocular system, our multi-camera setup enhances loop detection from different viewpoints, significantly increasing the detection rate of loops.
 TABLE I: Comparison of ATE on various datasets.
@@
 Trans. Rot. Scal. Trans. Rot. Scal. Trans. Rot. Scal. Trans. Rot. Scal. Trans. Trans. Trans. Trans.
-VINS-Fusion 76.198 114.923 1.003 38.213 125.079 0.585 14.749 123.580 0.181 135.536 130.248 2.924 16.579 10.921% 14.80 16.469%
-ORBSLAM3 22.673 117.981 0.390 12.269 115.832 0.306 8.134 114.361 0.164 44.317 114.268 1.484 2.90 1.91% 0.85 0.80%
-MultiCamSLAM 15.68(245.2) 115.516 0.341 8.91(222.5) 114.082 0.184 11.28(91.9) 155.944 0.150 56.46(553.5) 136.175 1.561 14.83 9.75% 1.69 1.88%
-MCVO(ORB) 44.054 8.634 0.518 64.861 4.803 1.367 49.143 5.353 0.591 102.050 6.139 2.833 11.50 7.561% 14.58 16.321%
-MCVO(ours) 13.933 6.847 0.300 20.234 2.585 0.394 6.895 3.341 0.079 76.282 7.348 2.638 7.894 5.200% 1.54 1.711%
+VINS-Fusion 76.188 30.918 1.008 38.213 37.502 0.556 14.516 30.054 0.181 39.235 33.082 1.464 16.579 10.921% 14.80 16.469%
+ORB-SLAM3 22.433 11.262 0.396 12.744 6.555 0.346 8.320 5.868 0.171 34.842 8.477 1.336 2.90 1.91% 0.85 0.80%
+MultiCamSLAM 15.307(245.2) 16.712 0.330 8.716(222.5) 10.382 0.181 6.744(91.9) 55.356 0.119 56.413(553.5) 43.253 1.568 14.83 9.75% 1.69 1.88%
+MCVO(ORB) 8.613 8.241 0.134 12.690 4.542 0.368 11.986 3.646 0.145 36.316 5.812 1.503 11.50 7.561% 14.58 16.321%
+MCVO(2-cam) 6.059 8.459 0.105 11.829 5.179 0.272 6.215 7.274 0.062 24.826 7.947 1.007 5.85 3.85% 4.35 4.83%
+MCVO(3-cam) 13.933 6.847 0.300 20.234 2.585 0.394 6.270 2.578 0.065 36.599 4.212 1.262 7.89 5.20% 1.54 1.711%
+MCVO(4-cam) 7.610 8.303 0.129 4.889 2.256 0.138 3.469 3.264 0.037 22.262 5.418 0.876 5.50 3.62% 0.95 1.06%
 Figure 5: Comparison of ATE for different methods on the KITTI360
-Pose Graph Optimization We follow the commonly used 4-DoF pose graph optimization [1] to update the trajectory in the loops.
-The pose graph $\bm{G}=\{\bm{V},\bm{E}\}$ keep track of all keyframes with their spatial constraints, where vertices $\bm{V}$ represent keyframes and edges $\bm{E}$ store the transformation constraints between keyframe poses.
-Keyframes marginalized during the sliding window are added to the pose graph, with their relative transformations to other keyframes in the local sliding window incorporated as edges.
-When a loop is detected, a loop edge between the current frame and the loop frame is added to $\bm{E}$, followed by the 4-DoF pose graph optimization.
-The optimization process effectively reduces accumulated errors and enhances the system’s robustness.
+Pose Graph Optimization We employ 6-DoF pose graph $\bm{G}=\{\bm{V},\bm{E}\}$ optimization [1] to refine trajectories using vertices $\bm{V}$ (keyframes) and edges $\bm{E}$ (transformations).
+Marginalized keyframes from the sliding window are incorporated into $\bm{G}$ with relative transformations as edges.
+Upon loop detection, a loop closure edge triggers 6-DoF optimization, mitigating accumulated errors and enhancing robustness through spatiotemporal constraint propagation.
 IV Experiments
-In this section, we conduct experiments on two public datasets to evaluate our proposed method.
 IV-A Experimental Setup
 IV-A1 Datasets
-For the evaluation of our method, the dataset must contain multi-camera setup with pre-calibrated intrinsic and extrinsic parameters.
-We select two public available dataset, i.e., KITTI360[5] and MultiCamData[7], which have different camera types, scenario types, and various camera layouts.
-The KITTI360 dataset was collected in outdoor road environments with two $180^{\circ}$ fisheye cameras and two $90^{\circ}$ pinhole cameras, as shown in Fig.
-1.
-The four cameras are synchronized and pre-calibrated.
-Compared to other datasets such as EuRoC [34], KITTI360 features larger scale, longer distances, and challenging scenarios, such as roads under bridges (lighting changes), wilderness (texture degradation), and dynamic scenes.
-We tested on most of the sequences and compared with the ground truth trajectories.
-MultiCamData was collected in indoor environments using 6 cameras.
-We mainly use the non-overlapping cameras (cam0, cam5, and cam6) for generalizability.
-The dataset also includes challenging scenarios, such as narrow corridors, large white walls (texture degradation), sudden turns, and dynamic scenes.
+For the evaluation of our method, we require datasets containing multi-camera setups with pre-calibrated intrinsic and extrinsic parameters.
+We select two publicly available benchmarks: KITTI-360 [5] and MultiCamData [7], which exhibit complementary characteristics in terms of camera configurations, environmental diversity, and operational challenges.
+The outdoor-oriented KITTI-360 dataset comprises 4 cameras with heterogeneous configurations: two $180^{\circ}$ fisheye and two $90^{\circ}$ pinhole cameras arranged on a vehicle (Fig.
+1).
+This dataset provides large-scale urban road scenarios with three principal challenges: (1) Illumination variations, (2) Texture degradation in wilderness, and (3) Dynamic objects.
+We evaluated on 4 sequences for rigorous trajectory comparison with ground truth.
+Besides, the indoor MultiCamData features 6 pinhole cameras comprising four forward and two side cameras.
+It presents distinctive challenges including (1) Feature scarcity in narrow corridors, and (2) Severe texture degradation on blank walls.
+The highly constrained environments enable effective verification of the robustness and generalization ability of the proposed method.
 IV-A2 Evaluation Setup
-We select VINS-Fusion [1], ORB-SLAM3 [2], and MultiCamSLAM [7] as baselines for comparison.
-To obtain metric scale state estimation for VINS-Fusion and ORB-SLAM3, these two methods are setup with front stereo cameras of each dataset.
-While for MultiCamSLAM, we employ a non-overlapping camera setup consistent with the method [7].
-MCVO is configured with two fisheye cameras and one front pinhole camera on KITTI360 dataset.
-The evaluation metrics on KITTI360 include Absolute Trajectory Error (ATE), Relative Pose Error (RPE) and scale error [35].
-Since MultiCamData lacks ground truth trajectory, we estimate the errors between the initial and final poses using a visual target as in [7], given that the start and end of the trajectories are at the same location.
-All the experiments are conducted on a desktop with a 2.5GHz Intel Core i7-11700 CPU and an NVIDIA RTX3060 GPU.
+We evaluate against three state-of-the-art visual odometry systems: VINS-Fusion [1], ORB-SLAM3 [2], and MultiCamSLAM [7].
+For metric consistency, VINS-Fusion and ORB-SLAM3 utilize calibrated stereo configurations, while MultiCamSLAM follows its native multi-camera implementation[10].
+Our method adopts four-camera fusion on KITTI-360: two fisheye + two pinhole cameras, with ablative configurations (2-cam: dual pinhole; 3-cam: dual fisheye + mono pinhole).
+Evaluation metrics comprise Absolute Trajectory Error (ATE), Relative Pose Error (RPE), and scale drift[33] for KITTI-360, while MultiCamData evaluate metrics via visual marker-based start-end pose discrepancy [7] due to the absence of ground truth trajectories.
+All experiments are conducted on a desktop with an Intel Core i7-11700 CPU and an NVIDIA RTX3060 GPU.
 IV-B Experimental Results
-The quantitative results for ATE are presented in Table I and Fig.
+The quantitative results for the ATE metric are shown in Table I and Fig.
 5.
-Compared with the stereo ORB-SLAM3 and VINS-Fusion, MCVO achieves significantly lower rotation error and competitive translation error on most of the sequences of KITTI360, primarily due to the wide FoV provided by the multi-camera setup and the high accuracy and robustness of the frontend.
-The small scale errors further demonstrate the effectiveness of our scale estimation strategy.
-The MultiCamSLAM method cannot finish most of the sequences on KITTI360.
-This issue arises because the KITTI360 dataset was captured at high speeds, leading to large inter-frame displacements.
-MultiCamSLAM struggles with significant inter-frame displacements, hindering triangulation and initialization.
-The comparison with MCVO on ORB features further shows the robustness of our front-end design.
-Besides, MCVO shows strong performance and robustness on MultiCamData.
-It outperforms VINS-Fusion in accuracy, with a marginal shortfall compared to ORB-SLAM3.
-This slight reduction in accuracy is primarily attributed to the presence of numerous white walls in indoor corridors, which introduce noise that impacts multi-camera VO.
+Stereo VINS-Fusion shows much worse results than stereo ORB-SLAM3.
+Compared with the stereo ORB-SLAM3 and VINS-Fusion, MCVO achieves lower rotation and translation error on most of the sequences of KITTI360.
+This improvement stems from two key advantages: (1) the robust scale estimation module reduces scale drift across all sequences, and (2) the multi-camera configuration enables stable feature tracking through wide FoV coverage.
+MultiCamSLAM fails to complete most sequences in KITTI360 due to inadequate handling of large inter-frame displacements in high-speed scenarios, preventing reliable triangulation and initialization.
+Experimental results of MCVO with varying camera configurations are presented in the last three rows of Table I.
+We first compare MCVO under 4 camera setup using ORB feature and the proposed learning-based front-end.
+The comparison shows the robustness of our front-end design.
+In terms of translation and scale estimation, 3-cam setup operates in a fully non-overlapping configuration, significantly outperforming ORB-SLAM3 and VINS-Fusion with stereo setup, which can be attributed to our robust multi-camera scale estimation strategy.
+Both 2-cam and 4-cam setups employ forward stereo cameras, where their overlapping FoV contributes to enhanced translation accuracy.
+Notably, the 4-cam setup generally achieves superior performance compared to the 2-cam setup, demonstrating the effectiveness of multi-camera configurations in improving state estimation accuracy.
+Regarding rotation estimation, multi-camera configurations significantly outperform stereo setups, as their distributed FoV maintains feature observability during rapid rotations, extending tracking continuity by 3-5 frames.
+Moreover, MCVO demonstrates consistent robustness on MultiCamData, notably outperforming VINS-Fusion in pose estimation accuracy while being competitive to ORB-SLAM3.
+This minor precision gap stems from texture-deprived environments like uniform white walls, that degrade feature matching fidelity.
 Compared to MultiCamSLAM, which also employs a non-overlapping camera setup, MCVO achieves superior accuracy and better generalization across different sequences.
 Fig.
-6 presents the RPE results of different methods on the 00 sequence of KITTI360 dataset.
-MCVO significantly outperforms other methods with smallest pose drift.
-This improvement is largely due to the wide coverage of scenes in the multi-camera setup with a robust and efficient frontend.
-The expanded FoV offers stronger constraints for state estimation, resulting in higher pose estimation accuracy.
-Additionally, the robust frontend ensures accurate feature tracking even in low-quality scenes.
+6 presents the RPE results of different methods on the 05 sequence of KITTI360.
+MCVO (4-cam) outperforms other methods with smallest pose drifts.
+This improvement is largely due to our continuous correction strategy and multi-camera wide FoV.
 These results validate the effectiveness of MCVO on state estimation.
-Figure 6: Comparison of RPE for different methods on the KITTI360_00 sequence.
-To further qualitative analyze the performance, we plot the trajectories of different methods on sequences 00 and 05 of KITTI360 dataset in Fig.
+To further qualitatively analyze the performance, we plot the trajectories of different methods on sequences 00 and 05 of KITTI360 in Fig.
 7.
 ORB-SLAM3 and VINS-Fusion exhibit large rotation errors on the 00 sequence.
-In contrast, MCVO shows improved stability and accuracy under such conditions.
-Additionally, the overall trajectory of our approach is smoother with more stable position estimation.
+In contrast, MCVO shows improved stability and accuracy under the same conditions.
+Figure 6: Comparison of RPE for different methods on the KITTI360_05 sequence.
 Figure 7: Trajectories of different sequences on KITTI360.
 IV-C Ablation Study
-In this section, we conduct ablation experiments for different numbers of cameras, feature extractors, feature matching, and whether to use loop closure.
-The effectiveness of each part of our proposed method is verified through comprehensive analyses.
-TABLE II: MCVO performance under different number of cameras on the KITTI360 00 sequence.
-Method Transl.[m] Rot.[°] UsageCPU[%]
-Mean Std Mean Std Mean
-2 cams 45.523 43.042 8.180 7.263 203.2
-3 cams 13.933 13.640 6.847 6.806 316.7
-4 cams 20.673 22.262 7.671 7.812 443.6
-IV-C1 Camera number
-Table II shows the state estimation accuracy of MCVO under different number of cameras.
-MCVO on two cameras obtains relatively lower accuracy due to the limited FoV, which provides insufficient information.
-When the number of cameras increases to 3, the estimation accuracy reaches its peak.
-However, with 4 cameras, there is a slight decrease in accuracy.
-This is primarily because two fisheye and one perspective cameras have already covered enough surrounding information, the added new camera will bring more observation noises and computational burden for the system.
-Consequently, the gain from the new camera cannot contribute to improving accuracy.
-TABLE III: MCVO performance under different feature extractors on the KITTI360 00 sequence.
-Method Transl.[m] Rot.[°] UsageCPU[%] FPS
-Mean Std Mean Std Mean
-Shi-Tomasi 39.390 40.612 9.256 8.290 398.3 18.75
-ORB 44.054 39.900 8.634 8.327 942.4 12.74
-SuperPoint 13.933 13.640 6.847 6.806 316.7 19.50
-IV-C2 Feature extractor
-We use pose accuracy and CPU usage metrics to compare different feature extraction methods.
-Table III shows that SuperPoint achieves the highest accuracy compared with commonly used Shi-Tomasi and ORB.
-This is mainly because SuperPoint can extract multi-level features and is more robust to environmental changes.
-Additionally, the CPU usage of MCVO(SuperPoint) is significantly lower.
-This reduction in CPU usage is because of our GPU-accelerated frontend.
-Consequently, The shift of CPU usage also ensures the computation resource for back-end optimization of accurate pose estimation and efficiency.
-IV-C3 Feature matching
-Table IV shows the performance of MCVO on different feature-matching methods.
-Flow-based methods exhibit lower CPU usage and more inliers compared to matching-based methods (KNN, SuperGlue).
-Additionally, we count the mean tracking times of each feature.
-LK shows the longest tracking times compared to others, which provides stable landmarks for the backend pose estimation.
-The small number of track times of other methods leads to the pose initialization failures.
-TABLE IV: MCVO performance under different feature matching methods (Superglue-SG).
+In this subsection, we conduct ablation study on the front-end and back-end of MCVO to verify their effectiveness.
+TABLE II: MCVO performance under different feature extractors on the KITTI360 00 sequence.
+Method Transl.[m] Rot.[°] UsageCPU[%] UsageGPU[%] FPS
+Mean Std Mean Std
+Shi-Tomasi(CPU) 9.413 9.448 8.570 8.559 534.68 - 18.75
+ORB(CPU) 8.613 8.595 8.241 8.217 942.4 - 12.74
+Superpoint(CPU) 28.841 26.653 9.553 9.108 1023.6 - 9.02
+SuperPoint(GPU) 7.610 7.793 8.303 8.445 495.42 43.75 19.50
+TABLE III: MCVO performance under different feature matching methods (Superglue-SG).
 Method Inlier UsageCPU Aver. track Initialization
@@
 SuperPoint+LK 99.61 66.41% 710.12% 16.40
+TABLE IV: MCVO performance under different feature selection on the KITTI360 00 sequence. (Spatial Feature Distribution-SFD)
+Method SFD $\downarrow$ Transl.[m] Rot.[°]
+Mean Std Mean Std
+score-based 24.25 8.740 8.678 8.617 8.515
+quad-tree 22.18 8.432 8.437 8.680 8.633
+3-priority 21.95 7.610 7.793 8.303 8.445
+IV-C1 Feature extractor
+Experimental evaluations comparing feature extractors across pose accuracy and computational efficiency metrics reveal SuperPoint’s superior performance (Table II).
+It achieves a 18.4% higher pose accuracy than conventional Shi-Tomasi and ORB, attributable to its multi-scale feature extraction capability and enhanced robustness to environmental variations.
+Furthermore, MCVO (SuperPoint) reduces CPU utilization by 47.4% compared to ORB, enabled by the GPU-accelerated frontend architecture.
+This efficiency gain facilitates resource allocation for backend optimization processes, ensuring both high-precision pose estimation and real-time operational efficiency.
+IV-C2 Feature matching
+As shown in Table III, feature matching based on LK flow in MCVO demonstrates superior computational efficiency and robustness compared to descriptor approaches.
+Quantitative analysis of feature tracking duration reveals that the LK flow maintains the longest feature lifespan (avg.
+16.18 frames), providing stable landmark constraints for backend optimization.
+In contrast, other methods exhibit short tracking lifespans, resulting in a high failure rate during pose initialization.
+IV-C3 Feature selection
+As shown in Table IV, we conducted ablation studies on the feature selection algorithm.
+We introduce a new metric, SFD, to quantify Spatial Feature Distribution by computing the feature density on uniform grids of images and derive the variance across all cells.
+Compared with score-based methods, our 3-priority algorithm demonstrates 9.48% higher feature dispersion, effectively alleviating the issue of excessive feature concentration in texture-rich regions.
+The proposed method also achieves 11.0% improvement in pose estimation accuracy over both quadtree and score-based algorithms, validating its effectiveness.
 IV-C4 Loop closure
-We also conduct experiments on the KITTI360 05 sequence to validate the effectiveness of the loop closure.
-The qualitative results are shown in Fig.
-8.
-We can observe that the multi-camera loop closure effectively performs loop detection and closure, providing stable global constraints for back-end optimization.
-MCVO successfully corrects the cumulative error in the map, achieving more precise trajectory estimation.
+To validate the efficacy of the multi-camera loop closure, we conducted systematic evaluations on the KITTI360 05 sequence.
+Qualitative analysis (Fig.
+8) demonstrates that our multi-camera closure framework enables robust loop detection and closure, establishing reliable global constraints for backend optimization.
+The proposed MCVO effectively mitigates cumulative mapping errors through this mechanism, achieving enhanced state estimation accuracy.
 Figure 8: Loop closure effectiveness validation on the KITTI360 05 sequence.
@@
   • Multi-camera configurations can improve VO robustness, but inevitably introduce new computational requirements.
-Our approach uses learning-based feature extraction to shift CPU consumption but introduces the need for GPU resources.
+Our approach uses learning-based feature extraction to shift CPU consumption but introduces the need of GPU resources.
 The best approach may be to develop a specialized camera with feature association embedded in the sensor hardware.
-  • MCVO requires the extrinsic parameters of arbitrarily arranged multi-cameras. When there is no overlap between FoV, the extrinsic calibration will be a challenging problem.
+  • MCVO requires the extrinsic parameters of arbitrarily arranged multi-cameras. When there is no FoV overlap between cameras, precise extrinsic calibration will also be a challenging problem.
 V Conclusions and Future Work
-In this paper, we propose a generic multi-camera visual odometry system, which has high flexibility on the camera setup with different camera types, with or without FoV overlap.
-The arbitrary arrangement of multi-cameras, not only improves the flexibility of the SLAM system but also achieves robust camera-only true-scale pose estimation.
-The designed learning-based feature association front-end utilizes GPU to effectively release the computational pressure of CPUs on multi-camera data processing and improves the stability of feature matching.
-The system functionalities of multi-camera features for scale initialization, back-end optimization, and loop closure are fully validated during the experiments.
-Comparisons with the state-of-the-art further demonstrate the effectiveness and robustness of MCVO on state estimation.
-Future work will focus on the deployment of MCVO on more platforms.
+In this paper, we propose a robust multi-camera visual odometry system, which has high flexibility on the camera setup with different camera types, with or without FoV overlap.
+The arbitrary arrangement of multi-cameras improves the flexibility of the SLAM system on more real-world platforms.
+Through establishing multi-camera manifold consistency constraints on motion trajectories, we effectively calculate metric-scale poses during initialization and continuously refine metric-scale in the backend optimization.
+Additionally, we further boost the accuracy and efficiency of the pose estimation through the learning-based feature association frontend and the multi-camera loop closure.
+Comparisons with the state-of-the-arts demonstrate the effectiveness and robustness of MCVO on state estimation.
+Future work will concentrate on deploying MCVO across more platforms and in diverse scenarios while simultaneously developing dense depth reconstruction using multi-cameras.